The combined area of a square and a rectangle is 132 square centimeters. The width of the rectangle is 2 centimeters more than the length of a side of the square, and the length of the rectangle is 4 centimeters more than its width. Find the dimensions of the square and the rectangle.
Answer by mananth(16946)

square side =x
area of square = x^2
width of rectangle = x+2
length of rectangle = x+2+4
area of rectangle = (x+2)(x+6)
(x+2)(x+6)+x^2= 132
x^2+8x+12+x^2=132
2x^2+8x+8\12-132=0
2x^2+8x-120=0
/2
x^2+4x-60=0
x^2+10x-6x-60=0
x(x+10)-6(x+10)=0
(x+10)(x-6)=0
x= 6
side of square = 6cm
width = 8cm
length =12 cm